OpenStudy (anonymous):

Linda was selling tickets for the school play. She sold 10 more tickets than children tickets and she sold twice as many senior tickets than children tickets. How do I solve for this?

OpenStudy (anonymous):

Well, you have a system of two equations. Call senior tickets s, and children's tickets c. Then: \[s = c+10\] \[s = 2c\] So, substitute 2c for s in the first equation: \[2c=c+10\] \[c = 10\] Plugging that back in to one of the original equations, s=20. There's your answer.